# Latent Diffusion Models for Controllable RNA Sequence Generation

### v1

@@ 4.1 Experimental Setup
4.2 Results: Samples from RNAdiffusion Resemble Natural ncRNAs
We report the test NLL of the latent generated distribution in Table 1. We see that increasing the query token length $K$ and the embedding dimension $D$ generally decreases the test NLL per dimension. This is expected from the information-theoretic view, as increasing the dimension of the latent space means that we will need fewer bits per dimension to represent the uncertainty sourced from the distribution of the discrete sequences. Nevertheless, doubling the total dimension of the latent space does not necessarily decrease the NLL by half, as this also allows the auto-encoder to retain more information in the latent space, which increases the entropy of the latent distribution.
For subsequent experiments, we choose to use the trained RNA auto-encoders that have average normalized edit distance $\leq 1\%$. They correspond to hyperparameters $(K,D)=(16,320)$, $(32,160)$, and $(32,320)$. We further select the parameter (32, 160) for generation, which achieves the smallest total NLL, i.e, NLL per dimension multiplied by the dimension of latent space ($K\times{D}$).
@@ 4.2 Results: Samples from RNAdiffusion Resemble Natural ncRNAs
First, generated sequences vary in length, with a distribution closely aligned with that of natural ncRNAs (see Figure 2). Second, we plot the distributions of the aforementioned metrics of our generated sequences and compare them to a held-out natural ncRNA reference set in Figure 3. Additionally, we create a curated random sequence set by replacing each token in the reference sequences with nucleotide randomly sampled at a probability of 0.5 for G/C, approximating the average G/C content ratio of natural ncRNAs, while maintaining the original sequence length.
The sequences generated by our latent diffusion model exhibit greater similarity to the natural sequences compared to the random sequences shown in Figure 3. Specifically, the Levenshtein and 4-mer distances underscore the proximity to the closest natural sequences. The results reveal that both generated and natural sequences exhibit similar modes in terms of distances to the nearest natural sequence whereas randomly generated samples display a distinct distribution. This implies that the generative model not only captures the contextual characteristics of natural RNA sequences but also produces novel sequences that maintain a minimum distance from natural samples. Additionally, generated sequences demonstrate a mean G/C content similar to that of natural samples while displaying a broader range of G/C content compared to random samples. The generated RNA sequences further exhibit distributions of Minimum Free Energies (MFEs) and secondary structures that considerably resemble natural patterns. This suggests that our latent diffusion model can generate stable RNA molecules with plausible secondary structures. Moreover, in the latent space shown in t-SNE maps, our generated sequences predominantly overlap with natural sequences and more accurately capture several separated modes presented in natural sequences. In contrast, random sequences fail to capture these distinct modes and display a larger shift from the natural distribution.
Finally, we compute the consensus secondary structures for various subsets of our generated RNA sequences and compared them against the natural sequences following [113]. Figure 6 demonstrates that the generated RNA sequences adopt secondary structures similar to those conserved in some common non-coding RNA families like tRNA, rRNA, and snRNA. Notably, our model does not explicitly condition on any specific families during generation.
@@ 5.2 Latent Reward Modeling for MRL and TE
Model Architecture. The latent reward networks take as input a list of $K$ latent vectors, obtained by the sequence autoencoder. We construct a 6-layer Convolutional Residual Network (ResNet) [53] specifically designed for 1D convolution, applicable to both the MRL and TE prediction tasks. Additional details are deferred to Appendix D.3.
Results. For each task, we train two reward models on top of the latent spaces from the pretrained RNA Sequence Autoencorder ($(K,D)=(32,160)$), each initialized under distinct random seeds. One model is utilized for guided generation in Section 5.3 while the other is employed for validation assessment. We calculate the Spearman and Pearson correlation coefficients between the predicted rewards and the actual rewards (shown in Table 2). Notably, the Spearman R values on the held-out test sets are $0.56$ for the TE task and $0.69$ for the MRL task, achieving slightly inferior performance compared with baselines [24, 103] and the state-of-the-art predictor built upon UTR-LM [30]. The results support our intuition that the embeddings after Q-Former remain useful for downstream tasks, allowing us to build effective reward models based on them.
Table 2: Performances of MRL and TE reward modeling. Numbers in parentheses show the agreement between the predictions of two reward models on the same dataset. Spearman R values of baseline models are from [30].
@@
Ours (agreement) UTR-LM [30] Cao et al. [24] Ours (agreement) UTR-LM [30] Sample et al. [103]
Test Spearman R 0.56 (0.90) 0.64 0.60 0.69 (0.79) 0.85 0.83
Test Pearson R 0.57 (0.91) - - 0.77 (0.88) - -
Train Spearman R 0.85 (0.95) - - 0.7547 (0.88) - -
Train Pearson R 0.86 (0.95) - - 0.8254 (0.90) - -
@@ 5.3 Reward-Guided Diffusion
(a) TE Optimization
(b) MRL Optimization
Figure 5: Performances of reward-guided diffusion on TE and MRL task. Baseline results refer to the evaluation of unguided generation with validation reward models. Guidance Reward Evaluation indicates the assessment with the same reward models used for guided generation. All evaluation sequences are over 30-bp.
We compute gradient-based forward universal guidance [17] in the latent space with the reward models trained in Section  5.2 to shift the latent backward sample paths towards generating high MRLs/TEs sequences; see Section 2.3. We choose the loss function $\ell$ to be $L_{2}$ loss and vary the target reward $r^{*}$ and the guidance strength $\lambda$. We take the latent reward models trained with one random seed for TE and MRL optimization. For cross-validation, we also employ the reward models trained with the other random seed as the oracle to evaluate the rewards of generated sequence.
Results. We tune the guidance by varying the value of guidance strengths $\lambda$ and target reward values $r^{*}$. Performances of output from our guided diffusion are shown in Figure 5. At a mild guidance strength, increasing the target TE or MRL values generally brings the reward improvements of generated sequences, demonstrating the efficacy of our reward-guided generation approach. Among all the guidance settings, The best validation TE ($1.84\pm 0.82$) is achieved with the moderate guidance strength as 800 and the largest log-target value as 5 , showing $166.7\%$ improvement compared to the validation TE of unguided generation ($0.69\pm 0.22$). Under the MRL-guidance setting of $\lambda=800$ and $\log r^{*}=4$, mean validation MRL of the generated sequences increases from $3.92\pm 1.31$ to $5.98\pm 0.90$ with maximum relative increase of $52.6\%$.
Guidance Trade-offs. For higher guidance strength, the improvements can be artificially amplified when evaluated with the same reward model used for guidance. Moreover, setting $\lambda$ and $r^{*}$ too large will eventually hurt the cross-validation performance, since the generation process is over-adapted towards the imperfect guidance reward model, which is a general phenomenon known as reward hacking [39, 32, 134].
To get insights into the phenomenon, we compare the generated distributions under different guidance strengths in terms of biological metrics (see Figure 9). With the $\lambda$ and $r^{*}$ increasing, the disparities in G/C content (lower mean) and Minimum Free Energy (higher mean) between the generated sequences and natural sequences become larger, which might be attributed to the distribution shift of the generated sequences. Adopting moderate $\lambda$’s and $r^{*}$’s would be important for reward-guided generation in terms of the trade-off between achieving higher reward values and maintaining generation qualities.
More concretely, we combine both reward evaluation and Minimum Free Energy (MFE) calculation into Pareto front curves (see Figure 8) and compare with two competitive baselines (see Section D.5 for details). Our gradient-guided latent diffusion model achieves higher rewards with the same MFE level compared to Random [103] and UTRGAN [18]. Although the sequences optimized by UTRGAN could achieve high rewards, they consistently exhibit high free energy. This demonstrates that our method has a better reward–structural stability trade-off compared to other baselines.
@@ 6 Related Work
People have been exploring building D3PM-type [14] discrete diffusion models and Diffusion-LM type [77] for proteins sequences and 3D structures [72, 127, 40, 6, 80, 121, 126, 131, 61, 45, 4, 88] and DNAs [15, 78, 37]. To the best of our knowledge, we are the first work to study building diffusion models for RNA sequences. Besides, latent diffusion models for text sequences have been studied in [128, 135, 87]. Due to space limits, we address the similarities and the differences of those related studies in Appendix C.3. For a detailed discussion on continuous/discrete diffusion models and their applications in biology, see the full related work section in Appendix C.
@@ 7 Conclusion
Our work presents the first application of latent diffusion models on RNA generation. Our pipeline involves a novel adaptation of the Querying Transformer to handle the inherently discrete, variable-length characteristics of RNA sequences. Through the incorporation of the gradient guidance of the latent reward models, our latent diffusion model can generate novel 5’-UTRs with $166.7\%$ higher mean TEs and $52.6\%$ higher mean MRLs. These results provide a biologically meaningful demonstration for RNA generation, a less explored yet exciting area of research.
@@ D.3 Additional Experimental Setup of Latent Reward Models
D.4 Additional Experimental Results for ncRNA Generation
(a) 5S ribosomal RNA
(b) U1 spliceosomal RNA
(c) tRNA
(d) 31 Generated Sequences
(e) 28 Generated Sequences
(f) 64 Generated Seqeunces
Figure 6: Consensus secondary structure diagrams of generated sequences compared to common non-coding RNA families following RfamGen [113].
D.5 Additional Experimental Results for Guided UTR Generation
Compared Baselines. We compare our TE/MRL-guided generation method with two competitive baselines and plot the Pareto front curves between MFE and TE/MRL values (see Figure 8).
  • Random. We adopt a best-of-$N$ sequence generation method similar to Sample et al. [103], which uses a genetic algorithm to generate random sequences and merely retain ones with high rewards. In our approach, we generate $N$ random sequences for each natural sequence in the same way shown in Section  4.2 and select the sequence with the highest reward according to our guidance reward model.
  • UTRGAN, a novel 5’-UTR sequence generative model targeting translation rate optimization. We directly generate 6000 reward-optimized sequences using the official codes release by UTRGAN [18].
(a) MRL
(b) TE
Figure 8: Pareto front curves between Minimum Free Energy (MFE) and Mean Ribosome Loading (MRL)/ Translation Efficiency (TE). The curves are generated by selecting the top 10% quantile MRL/TE within sliding windows around each MFE value. Each shaded dot denotes a sequence.
(a) G/C Content
(b) Minimum Free Energy
D.6 Ablations Studies
We conducted ablations on RNA-FM and Q-former in Table 10(b) & 4. The results demonstrate the necessity of main components of our design:
  • When we replace the pretrained RNA-FM with naive one-hot embeddings, both the reconstruction capability of the sequence autoencoder and the reward model performance are affected, where we observe significant increases in the reconstruction error and declines in the R-values of the reward model (Table 10(b)).
  • When we replace the Q-former with a naive average pooling, the reconstruction performance of the Sequence AutoEncoder is significantly dampened, resulting in a 34.8% normalized edit distance (Table 4).
Table 3: Ablations on RNA-FM. We replace RNA-FM with naive one-hot token embedding and test the performances on sequence autoencoder and MRL reward model training under different Q-former settings ($K$, $D$). NLL: Reconstruction Error.
@@ D.6 Ablations Studies
NLL $\downarrow$ Normalized Edit Distance $\downarrow$ Test Spearman R $\uparrow$ Test Pearson R $\uparrow$
@@
NLL $\downarrow$ Normalized Edit Distance $\downarrow$ Test Spearman R $\uparrow$ Test Pearson R $\uparrow$
@@
Settings ($K\times{D}$) Sequence AutoEncoder
NLL $\downarrow$ Normalized Edit Distance $\downarrow$
w/o Q-former ($1\times 640$) 1.0061 34.78 % $\pm$ 17.81 %
w/ Q-former ($16\times 160$) 0.0156 2.19% $\pm$ 5.50 %

### v2

@@ 4.1 Experimental Setup
4.2 Results: Samples from RNAdiffusion Resemble Natural ncRNAs
We report the test NLL of the latent generated distribution in Table 1. We see that increasing the query token length $K$ and the embedding dimension $D$ generally decreases the test NLL per dimension. For subsequent experiments, we choose to use the trained RNA auto-encoders that have average normalized edit distance $\leq 1\%$. They correspond to hyperparameters $(K,D)=(16,320)$, $(32,160)$, and $(32,320)$. We further select the parameter (32, 160) for generation, which achieves the smallest total NLL, i.e, NLL per dimension multiplied by the dimension of latent space ($K\times{D}$).
We compare the generated samples (with $K=32$, $D=160$) with natural ncRNAs. We plot the length distribution of the sequences in Figure 2 and demonstrate that the generated sequences vary in length, with a distribution closely aligned with that of natural ncRNAs.
@@ 4.2 Results: Samples from RNAdiffusion Resemble Natural ncRNAs
Additionally, we create a curated random sequence set by replacing each token in the reference sequences with nucleotide randomly sampled at a probability of 0.5 for G/C, approximating the average G/C content ratio of natural ncRNAs, while maintaining the original sequence length. Next, we plot the distributions of the aforementioned metrics of our generated sequences and compare them to a held-out natural ncRNA reference set and the set of random sequences in Figure 3, which demonstrates that our generated sequences exhibit greater similarity to the natural sequences compared to the random sequences in terms of sequence-level metrics (Levenshtein and 4-mer distances, G/C content ratios), structural-level metrics (Minimum Free Energies and secondary structure distances), and latent space similarities (t-SNE maps).
@@ 5.2 Latent Reward Modeling for MRL and TE
Model Architecture. The latent reward networks take as input a list of $K$ latent vectors obtained by Q-former. We construct a 6-layer 1D convolutional neural network with residual blocks [53] and apply it on both the MRL and TE prediction tasks. Additional details are deferred to Appendix D.3.
Results. For each task, we train two reward models on top of the latent spaces from the pretrained RNA Sequence Autoencorder ($(K,D)=(32,160)$), each initialized under distinct random seeds. One model is utilized for guided generation in Section 5.3 while the other is employed for validation assessment. We calculate the Spearman correlation coefficients between the predicted rewards and the actual rewards (shown in Table 3). Notably, the Spearman R values on the held-out test sets are $0.56$ for the TE task and $0.69$ for the MRL task, achieving slightly inferior yet comparable performance compared with baselines [24, 101]) and the state-of-the-art predictor built upon UTR-LM ( [30]).
Table 3: Spearman R-values of MRL and TE reward modeling. The embeddings after Q-Former remain useful for downstream tasks, allowing us to build effective reward models based on them.
@@
Ours UTR-LM [30] [24] Ours UTR-LM [30] [101]
Test Spearman R 0.56 0.64 0.60 0.69 0.85 0.83
@@ 5.3 Reward-Guided Diffusion
We compute gradient-based forward universal guidance [17] in the latent space with the trained reward models to shift the latent backward sample paths towards generating high MRLs/TEs sequences as described in Section 2.3, where we choose the loss function $\ell$ to be $L_{2}$ loss and vary the target reward $r^{*}$ and the guidance strength $\lambda$. For guidance, we take the latent reward models trained with one random seed for TE and MRL optimization. For cross-validation, we employ the reward models trained with another random seed as the validation model to evaluate the rewards of the generated sequence. The best performing guidance parameters are able to improve the mean TE by up to $166.7\%$ and mean MRL by up to $52.6\%$ compared to the unguided baseline. We defer our systematic analysis of the effects of different guidance parameters to Appendix D.6.
Baselines. We compare the performance of our guidance pipeline with the following two baselines:
  • Random. We adopt a best-of-$N$ sequence generation method similar to Sample et al. [101], which uses a genetic algorithm to generate random sequences and merely retain ones with high rewards. In our approach, we generate random sequences as in Section  4.2 and select the sequence with the highest reward according to our guidance reward model.
  • UTRGAN, a 5’-UTR sequence generative model targeting translation rate optimization. We directly generate 6000 reward-optimized sequences using the official codes release by UTRGAN [18].
Results. For each generated sequence, we evaluate the reward using the validation reward models and also calculate the Minimum Free Energy (MFE) as the proxy for structural stability. We plot the validation reward (MRL/TE) v.s. MFE of samples generated by different methods in the Pareto fronts (Figure 5). Our gradient-guided latent diffusion model achieves higher rewards with the same MFE level compared to the two baselines. This demonstrates that our method has a better reward–structural stability trade-off.
(a) MRL
(b) TE
Figure 5: Pareto front curves between Minimum Free Energy (MFE) and Mean Ribosome Loading (MRL)/Translation Efficiency (TE). The curves are generated by selecting the top 10% quantile MRL/TE within sliding windows around each MFE value. Each shaded dot denotes a sequence. All the sequences are evaluated with the same validation reward models.
@@ 6 Related Work
People have been exploring building D3PM-type [14] discrete diffusion models and Diffusion-LM type [77] for proteins sequences and 3D structures [72, 123, 40, 6, 80, 117, 122, 127, 61, 45, 4, 88] and DNAs [15, 78, 37]. Besides, latent diffusion models for text sequences have been studied in [124, 131, 87]. For a detailed discussion on continuous/discrete diffusion models and their applications in biology, see the full related work section in Appendix C.
@@ 7 Conclusion
We present RNAdiffusion , a diffusion-based generative model that captures the RNA sequence distribution within a latent space constructed by a Sequence AutoEncoder, which utilizes a Querying Transformer to manage the variable-length nature of RNA sequences. By incorporating gradient guidance from latent reward models, our diffusion model can generate novel 5’-UTRs with up to $166.7\%$ higher TE and $52.6\%$ higher MRL, while achieving a superior trade-off between rewards and structural stability. These results provide a biologically relevant demonstration of RNA sequence generation, an underexplored yet exciting area of research.
Acknowledgments
We thank Zaixi Zhang for valuable suggestions to the paper.
@@ D.1 Additional Experimental Setup of RNA AutoEncoder
Details of the Models of Sequence AutoEncoder. We adopt the pretrained RNA-FM [26] as the encoder. RNA-FM is an encoder-only model trained with masked language modelling on non-coding RNAs (ncRNAs), which can extract biologically meaningful embeddings from ncRNAs and can be adapted for various downstream tasks, such as UTR function prediction, RNA-protein interaction prediction, secondary structure prediction, and etc. For the Q-Former, we modify the ESM-2 model architecture [82] by adding $K$ trainable query token embeddings and adding cross-attention layers in each transformer block that attends to the embeddings from the RNA-FM. For the decoder model, we adopt the same model architecture of ProGen2-small [93] with our customized tokenizers for RNAs.
@@ D.3 Additional Experimental Setup of Latent Reward Models
D.4 Ablations Studies on RNA-FM
We validate the necessity of the pre-trained RNA-FM in this section. When we replace the pre-trained encoder RNA-FM with naive one-hot embeddings, both the reconstruction capability of the sequence autoencoder and the reward model performance are affected, where we observe significant increases in the reconstruction error and declines in the R-values of the reward model (Table 6(b)).
Table 4: Ablations on RNA-FM. We replace RNA-FM with naive one-hot token embedding and test the performances on sequence autoencoder and MRL reward model training under different Q-former settings ($K$, $D$). NLL: reconstruction error, NED: normalized edit distance.
@@ D.4 Ablations Studies on RNA-FM
NLL $\downarrow$ NED $\downarrow$ Test Spearman R $\uparrow$ Test Pearson R $\uparrow$
@@
NLL $\downarrow$ NED $\downarrow$ Test Spearman R $\uparrow$ Test Pearson R $\uparrow$
@@
D.5 Evaluation of 5’-UTR RNAdiffusion Fine-tuning
We compare the samples generated by fine-tuned latent diffusion models with natural 5’UTR reference dataset and random samples following the same evaluation procedure in Section 4.1. The results are shown in Figure 6.
D.6 Ablations on Guidance Parameters for Guided 5’UTR Generation
Guidance Parameter Search. We tune the guidance for generation by varying the values of guidance strength $\lambda$ and target reward $r^{*}$ as of Equation 5, and plot the rewards of guided generation under different guidance parameters $\lambda$ and $r^{*}$ in Figure 7. Among all the guidance settings, the best validation TE ($1.84\pm 0.82$) is achieved with the moderate guidance strength as 800 and the log-target value as 5 , showing 166.7% improvement compared to the validation TE of unguided generation ($0.69\pm 0.22$). Under the MRL-guidance setting of $\lambda=800$ and $\log r^{*}=4$, mean validation MRL of the generated sequences increases from $3.92\pm 1.31$ to $5.98\pm 0.90$ with maximum relative increase of 52.6 %.
(a) TE Optimization
(b) MRL Optimization
Figure 7: Performances of reward-guided diffusion on TE and MRL tasks. Baseline results refer to the evaluation of unguided generation with validation reward models. Guidance Reward Evaluation indicates the assessment with the same reward models which are used for guided generation. All evaluated sequences are over 30-bp.
Biological Metrics. We compare the generated distributions under different guidance strengths in terms of biological metrics (see Figure 8). With the $\lambda$ and $r^{*}$ increasing, the disparities in G/C content and Minimum Free Energy (MFE) between the generated sequences and natural sequences become larger. Specifically, the distribution of G/C content is shifting downwards while the mean MFE is increasing, indicating that the stability of generated sequences has been affected.
Reward Hacking in Reward-guided Diffusions. As shown in Figure 7, at a mild guidance strength, increasing the target TE or MRL values generally brings the reward improvements of generated sequences, demonstrating the efficacy of our reward-guided generation approach. For higher guidance strength, the improvements can be artificially amplified when evaluated with the same reward model used for guidance. Continuing to increase $\lambda$ and $r^{*}$ will eventually hurt the cross-validation performance, since the generation process is over-adapted towards the imperfect guidance reward model, which is a general phenomenon known as reward hacking [39, 32, 130]. Additionally, increasing $\lambda$ or $r^{*}$ may cause the distribution shift of biological properties of sequences such as G/C content and MFE (see Figure 8), affecting the stability of UTR sequences. Thus, adopting moderate $\lambda$’s and $r^{*}$’s would be important for reward-guided generation in terms of the trade-off between achieving higher reward values and maintaining generation qualities.
(a) G/C Content
(b) Minimum Free Energy